OpenStudy (anonymous):

Find the difference. (Difference of Rational Expressions) WILL GIVE MEDAL AND WILL BECOME FAN. NEED HELP QUICK! 5/x^2-6x - 1/x^2-36

OpenStudy (anonymous):

\[\frac{ 5 }{ x^2-6x }- \frac{ 1 }{ x^2-36 }\]

OpenStudy (anonymous):

is it 360

OpenStudy (anonymous):

says its incorrect

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

i tried my best

OpenStudy (anonymous):

\[\frac{ 4 }{ x^2+4x }- \frac{ 7 }{ x^2-16 }\]

OpenStudy (anonymous):

& thank you though (:

OpenStudy (anonymous):

you think you might know this one?

OpenStudy (anonymous):

your welcome and yes i do

OpenStudy (anonymous):

-12 im not sure this is correct wait one more min

OpenStudy (anonymous):

figure it out yet? (:

OpenStudy (anonymous):

@Nnesha think you cold help? someone said to tag you

OpenStudy (anonymous):

@paki